Find the differential equations of the space curve in which the two families of surfaces

u = x2 - y2 = c1 and v = y2 - z2 = c2 intersect.


1
Expert's answer
2021-06-01T03:41:09-0400

Ans:-

Given two families of surfaces

"u=x^2-y^2=c_1 \ \ \ \ and \ \ \ v=y^2-z^2= c_2"

For all level surface

"du=0 \to du=2xdx-2ydy=0 \ \to\\
\ \ \ \ \ \ \ \ \ \ \ \ xdx=ydy"


"\ \ \ \ dv=0 \to dv=2ydy-2zdz=0 \ \ \ \to \\
 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ ydy=zdz"

Then,

"xdx=ydy=zdz\u2223\u00f7(xyz)"


"\dfrac{xdx}{xyz}=\dfrac{ydy}{xyz}=\dfrac{zdz}{xyz}"


"\dfrac{dx}{yz}=\dfrac{dy}{xz}=\dfrac{dz}{xy}" is auxiliary equation


Conclusion

"(yz)p+(xz)q=xy" is desired equation